Find the slope of the line through the points (4, 7) and (-6, 2)
Answer by TimothyLamb(4379)

Calculate slope: m = (y2 - y1)/(x2 - x1)
---
p1 = (4, 7)
p2 = (-6, 2)
---
m = (2 - 7)/(-6 - 4)
m = -5 / -10
m = 1/2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php